OpenStudy (anonymous):

What is the approximate circumference of a circle with a radius of 30 inches? Use π ≈ 3.14. A. 62.8 inches B. 94.2 inches C. 47.1 inches D. 188.4 inches

OpenStudy (anonymous):

@igreen

OpenStudy (jordan123321):

what do you think

OpenStudy (anonymous):

idk

OpenStudy (jordan123321):

just try

OpenStudy (texaschic101):

there is a formula for this... C = 2 * pi * r

OpenStudy (jordan123321):

b or d which one

OpenStudy (anonymous):

how do i find the answer. so i use the formula then what? what does the formula mean. could you type the steps?

OpenStudy (texaschic101):

formula for circumference of a circle.... C = 2 * pi * r C = 2 * 3.14 * 30...now just do the math and multiply everything

OpenStudy (anonymous):

d

OpenStudy (jordan123321):

tht is right

OpenStudy (texaschic101):

correct...but if you were given the diameter instead of the radius, you would use this formula : C = pi * r

OpenStudy (anonymous):

thanks so much

OpenStudy (texaschic101):

knowing these 2 formulas : C = 2 * pi * r C = pi * d whether your given radius or diameter, you will be able to find the circumference of a circle